OpenStudy (anonymous):

Which of the following is the correct graph of the solution to the inequality −8 −5x + 2 > −38?

OpenStudy (aripotta):

can you show the graphs?

OpenStudy (anonymous):

http://gyazo.com/ce970df6fb62828549eda3f44292111c

OpenStudy (anonymous):

You need to solve it first.

OpenStudy (anonymous):

There is an inequality missing I think: −8 (?) −5x + 2 > −38? Is it\[\le or \ge \]

OpenStudy (anonymous):

Greater than or equal to >= Less than or equal to <=

OpenStudy (anonymous):

−8 ≥ −5x + 2 > −38?

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok so you need to solve each inequality: -8 ≥ -5x + 2 -10 ≥ -5x Remember to flip the sign when you are dividing by a negative\[2 \le x\]\[x \ge 2\] −5x + 2 > −38 -5x > -40 Remember to flip the sign when you are dividing by a negative x < 8 So x ≥ 2 and x < 8 Greater than... shading to the right Less than...shading to the left Equal sign... Closed circle No equal sign...Open circle

OpenStudy (anonymous):

Can you graph it now?